OpenStudy (anonymous):

If secθ=−5/4with π/2<θ<π , find the exact value of the other 5 trig functions of θ . If secθ=−5/4with π/2<θ<π , find the exact value of the other 5 trig functions of θ . @Mathematics

OpenStudy (mertsj):

sin = 3/5, cos = -4/5, tan = -3/4 cot = -4/3 csc = 5/3

OpenStudy (anonymous):

think of a 3 - 4 - 5 right triangle, take the appropriate ratios and be careful of the signs.

OpenStudy (anonymous):

why is tan negative?